Assume that today is December 31, 2019, and that the following information applies to Abner Airlines: After-tax operating income [EBIT(1 - T)] for 2020 is expected to be $700 million. The depreciation expense for 2020 is expected to be $180 million. The capital expenditures for 2020 are expected to be $350 million. No change is expected in net operating working capital. The free cash flow is expected to grow at a constant rate of 5% per year. The required return on equity is 15%. The WACC is 10%. The firm has $208 million of non-operating assets. The market value of the company's debt is $4.742 billion. 330 million shares of stock are outstanding. Using the corporate valuation model approach, what should be the company's stock price today? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Free Cash Flow (FCF) = EBIT x (1 - T) + Depreciation - NWC - Capex

= 700 + 180 - 350 = 530 million

EV = FCF / (WACC - g) = 530 / (10% - 5%) = 10,600 million

EV = Equity + Debt - Non operating assets

=> Equity Value = 10,600 + 208 - 4,742 = 6,066 million

Share Price = 6,066 / 330 = $18.38